OpenStudy (anonymous):

A 20.0 N bucket is to be raised at a constant speed of 50 {\rm cm/s} by a rope. According to the information in the table (approximate breaking strengths), how many kilograms of cement can be put into this bucket without breaking the rope if it is made of Thin white string 50 N ..

OpenStudy (anonymous):

This is the same as your earlier problem. Since there is no acceleration, the force of tension in the wire will be equal to the weight of the bucket/cement combination. Therefore, if we make the wire from thin white string, we can add 30 N of cement. Since \(50 = 20 + x\)

OpenStudy (anonymous):

k ty

OpenStudy (anonymous):

when there is a force applied ona body does it cause it to accelerate??

OpenStudy (anonymous):

If the net external forces are greater than zero, yes. This is Newton's Second Law. Note that here the bucket is travelling at CONSTANT velocity. Therefore, there is no acceleration since, \[a = {\Delta v \over \Delta t} = {0 \over \Delta t} = 0\] Let's draw a FBD about our bucket. If there is no acceleration, then we know from Newton's Second Law that\[F = m \cdot a = 0\]Therefore, \[T - mg = 0\]